Please, how do i express this as an equation:
The total attendence for two successive nights was 350. The second night, 50 more people attended than the first night. How many attended each night?
I tried: x+y=350
y=50+x (not sure on this part)

Found 3 solutions by jim_thompson5910, LisaJ, Alan3354:
Answer by jim_thompson5910(28717)

You have the correct equations. Nice work.

Since y = 50+x, we can replace 'y' in the first equation with '50+x' to go from

x+y=350

to

x + 50+x = 350


From here, solve for x. Then use this solution for x to find the solution for y.

Answer by LisaJ(11)

Yes, your two equations are OK.
Do you know how to solve the 2 equations together?

Answer by Alan3354(31538)

The total attendence for two successive nights was 350. The second night, 50 more people attended than the first night. How many attended each night?
--------
a = 1st night's attendance.
a + 50 = 2nd night's
------
a + a+50 = 350
a = 150
2nd night = 200